OpenStudy (logan13):

A cylinder has a radius of 14 m and a height of 6 m. What is the exact volume of the cylinder?

OpenStudy (logan13):

http://prntscr.com/aasqw3

OpenStudy (isaidavila):

V=πr^2*h

OpenStudy (logan13):

D

OpenStudy (logan13):

@rishavraj

OpenStudy (logan13):

@KyanTheDoodle

OpenStudy (logan13):

@welshfella

OpenStudy (kyanthedoodle):

Try doing the math without timing it by pi. Whatever number you get should be one of the answers.

OpenStudy (michele_laino):

hint: we have to apply the subsequent formula: \[\huge V = \pi {r^2}h = \pi \times 14 \times 14 \times 6 = ...?\]

OpenStudy (logan13):

504

OpenStudy (michele_laino):

I'm sorry I think it is wrong

OpenStudy (logan13):

wait i put in wrong

OpenStudy (logan13):

c

OpenStudy (michele_laino):

hint: what is \(14 \times 14 \times 6 =...?\)

OpenStudy (logan13):

A